### v5

@@ 6.2.1. Implications for practice and policy
For practitioners and policy makers, Tables 5 - 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection. These tables also reveal gaps that are relevant for conformity assessment under emerging regimes. Applied teams can use the matrix to assemble a minimal set of evaluations that cover privacy, robustness, transparency, logging, and post-market monitoring, and then plan additions for multimodal deepfake risk and adaptive red teaming as required by sectoral context.
6.3. Evidence portability and benchmarking
6.3.1. Limitations and threats to validity
This survey follows a pre-specified protocol, focusing on the period from November 2022 to August 2025, and incorporates selected foundational works to ensure definitional clarity. Nevertheless, decision regarding selection and synthesis may introduce bias. Reliance of public papers, dataset cards, and documentation implies that certain capabilities may be under-reported or reported with delay. The rubric weights are assigned equally by default and criterion coverage is determined through structured reading of sources, followed by discussion and adjudication in case of disagreements. Replication of our screening and scoring is facilitated by the availability of the documented sources and explicit inclusion rules.
6.4. Research directions.
@@ 6.3.1. Limitations and threats to validity
We highlight five priorities. First, extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps). Second, fold privacy and provenance checks into routine testing, including membership-inference and training-data extraction probes, plus audits of logging sufficiency. Third, create deepfake risk assessments for multimodal models that cover detection, content provenance, and robustness to evasion. Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and tests under distribution shift. Fifth, report sustainability alongside accuracy and robustness—energy, emissions, and compute-normalized efficiency—to enable compute-proportional evaluation and responsible deployment. These priorities directly address the risk-surface gaps surfaced in this survey.
6.4.1. Guidance for practitioners
@@
Teams putting responsible GenAI into practice can follow a simple loop. First, scope risks by mapping planned features to common risk families (e.g., bias, toxicity, privacy, and security). Next, select baselines from our matrix so that each release is covered by minimum tests across those areas. Then, add context: turn on logging, enforce dataset/model versioning, and keep lightweight documentation so decisions and results remain traceable. Before launch—and later during post-market monitoring—exercise adaptivity with red teaming and a few undisclosed checks to catch regressions. Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, inclusivity indicators, and energy/emissions. This routine keeps evaluation aligned with governance expectations and makes evidence easier to reuse when audits occur.
@@ 7. Conclusion
This survey examined responsible GenAI through the lens of governance needs and practical evaluation. We organized the literature into risk families and assurance qualities and introduced a two block rubric that distinguishes risk coverage (C1 to C8) from assurance and governance qualities (C9 to C10). From this review article, three findings stand out. First, coverage is strongest for bias and toxicity and weaker for privacy, deepfakes and provenance, and system level failure in tool using or agentic settings. Second, most evaluations remain static and task local, which limits evidence portability for governance and audit. Third, documentation and metric validity vary substantially, which complicates interpretation across releases and domains. These gaps are significant because contemporary systems are multimodal, often connected to tools and data stores, and operate under distribution shift.
We outlined a research agenda that calls for evaluations addresing system level failure, privacy leakage and provenance, deepfake and media integrity risk, and adaptive adversaries. We further recommended stronger reporting of constructs, metrics, uncertainty, and versioned artifacts, to enhance the portability of the results across audits and over time. For practitioners, we provided a sequence that links product risks to a minimal set of tests, adds assurance artifacts such as documentation and logging, and incorporates red teaming and post release monitoring. Collectively, these steps align engineering practice with governance expectations without prescribing a single benchmark.
This review has several limitations. It focuses on the period from November 2022 through August 2025 with selective inclusion of foundational work, relies on public documentation, and uses equal weights by default for the composite index. We mitigated subjectivity through a prespecified protocol, independent scoring, adjudication, and the release of a codebook, yet some capabilities may be under reported or reported with delay. The governance crosswalk is intended as plausible evidentiary support rather than a statement of legal compliance. Responsible GenAI requires progress on methods and on evidence. Benchmarks should move beyond static probes toward evaluations that reflect real operating conditions, including adaptivity, provenance, and end to end behavior. Governance frameworks should in turn recognize evidence that is documented, reproducible, and linked to explicit constructs. By connecting risk families to concrete tests and by foregrounding assurance qualities, this survey provides a path for researchers and practitioners to close capability gaps, to increase trust, and to support safe and accountable deployment of GenAI systems across domains.
@@
This quote encapsulates the core challenge of designing AI systems that can adapt ethically and responsibly to evolving societal norms. Adaptability in AI must go beyond technical performance to include ethical flexibility, ensuring systems align with human values. RAI emphasizes not just technical capabilities but also the ethical imperatives of transparency, fairness, and accountability.
@@
Achieving RAI requires a collective effort among developers, policymakers, and users to mitigate bias, prioritize privacy, and ensure accountability. Collaboration across disciplines and sectors is essential to create equitable and ethical AI systems that serve diverse societal needs.
@@
Traditional AI benchmarks have prioritized accuracy and efficiency, often neglecting fairness, accountability, and societal impact. Redefining benchmarks to incorporate ethical metrics is critical for guiding the development of AI systems that perform both technically and ethically. Such benchmarks must evaluate transparency, safety, and bias mitigation alongside traditional performance metrics.
The data-driven nature of AI presents persistent challenges, as models often inherit biases and privacy risks from training data. These biases, embedded in AI outputs, require rigorous mitigation strategies to distinguish between useful correlations and harmful patterns. While creating a perfectly unbiased AI system is unattainable due to the evolving nature of societal values (Ai et al., 2018), continuous evaluation and adaptation of AI models are essential to maintain alignment with ethical standards.
Human decision-making is influenced by moral reasoning and context, while AI operates strictly within its programmed algorithms. This underscores the importance of fostering AI literacy (Long and Magerko, 2020) among developers and users to ensure responsible deployment and use. To implement RAI effectively, AI systems must prioritize transparency, fairness, and adaptability. They should incorporate ethical reasoning and evolve alongside societal norms. As we advance toward Artificial General Intelligence, the true measure of AI should extend beyond technical proficiency to include ethical adaptability (Blili-Hamelin et al., 2024).
The pursuit of RAI is challenging yet indispensable. Ongoing advancements in RAI frameworks, coupled with a growing emphasis on AI ethics, offer a path forward. Prominent voices in the field, including Jensen Huang and Sam Altman, emphasize the need for safe and ethical AI. By aligning technical innovation with ethical principles, we can build AI systems that uphold accountability, foster trust, and adapt responsibly to the complexities of a dynamic world.
@@ 8. Acknowledgement
Appendix
Appendix A Full Search Strings and Filters
To enable exact replication, we report verbatim queries and filters for each source. For all databases, we limited the date range to November 2022–August 2025 and restricted results to English. Field constraints (title/abstract/keywords) were used where supported; otherwise, queries were applied to full text. Below we list the precise strings and any database-specific options.
ACM Digital Library.
Fields: Title/Abstract/Keywords. Years: 2022–2025. Document types: Journal, Proceeding Article.
@@ Appendix A Full Search Strings and Filters
IEEE Xplore.
Metadata search (Abstract + Index Terms). Years: 2022–2025. Content type: Journals, Conferences.
@@
SpringerLink / ScienceDirect.
Title/Abstract/Keywords. Years: 2022–2025. Subject areas: Computer Science, AI.
("Responsible AI" OR "AI governance" OR "AI safety" OR fairness OR transparency
 OR accountability OR privacy) AND ("generative AI" OR LLM OR LVLM OR multimodal)
Web of Science.
Topic search (TS=Title, Abstract, Keywords). Years: 2022–2025.
@@
DBLP.
Title search; manual verification of scope.
("Responsible AI" OR "Ethical AI" OR "AI safety" OR fairness OR transparency
 OR accountability OR privacy) AND (LLM OR LVLM OR "generative AI" OR multimodal)
arXiv.
Categories: cs.AI, cs.LG, cs.CV, cs.CL. Date filter applied via API.
@@
Google Scholar.
Custom range: Nov 2022–Aug 2025; “include citations” disabled during screening.
"Responsible AI" OR "AI governance" OR "AI safety" fairness transparency accountability
 privacy "generative AI" LLM LVLM multimodal
For each included study we recorded venue and year, dominant modality, problem area, method type, risks and controls addressed, datasets or artifacts, evaluation facets and metrics, and the type of evidence advanced. We adapted appraisal criteria from established CS/SE review checklists to rate method transparency, data/source adequacy, discussion of validity threats, and presence of reproducibility signals. Each study received qualitative notes and an ordinal rating (Low/Medium/High) per criterion; we also assigned an evidence-strength tag that informed interpretation during synthesis.
Figure A.1. Taxonomy diagram categorizing the core concepts of Responsible AI (RAI) into principal subsets: Explainable AI (XAI), Ethical Considerations, Sector Applications, Advanced Models, Privacy and Security, and Regulatory Frameworks. Each subset is further dissected into detailed elements such as Transparency, Bias Mitigation, Healthcare, Generative Models, and Data Anonymization, illustrating the flow from general themes to specific methodologies within RAI:
Figure A.2. Framework for Responsible Generative AI emphasizing transparency, fairness, accountability, privacy, and bias mitigation through the integration of data, models, regulations, and user-centric practices
Appendix B Codebook for rubric C1 to C10
Use. Score only explicit tasks and metrics. If a task maps to multiple risks, assign it to the primary stated objective to avoid double counting.
Evidence priority. Official paper and docs; then dataset or model cards and maintained leaderboards; then widely cited third party reports.
@@ Appendix B Codebook for rubric C1 to C10
Appendix C Scoring protocol, reliability, and sensitivity
Procedure. Two reviewers score each criterion using this codebook; disagreements resolved by a third reviewer. Publish the codebook, score sheets, and evidence extracts.
Reliability. Report Cohen’s $\kappa$ per criterion and the mean $\bar{\kappa}$; target $\bar{\kappa}\geq 0.60$.
Indices. Report the risk coverage index $S_{\mathrm{cov}}$, coverage rate $R$, assurance index $S_{\mathrm{assur}}$, and the overall normalized index $S$ as defined in Section 4.2.1.
Sensitivity. Recompute $S$ under alternate weights that prioritize governance critical risks and report Spearman rank correlation with the equal weight baseline.
@@ Appendix C Scoring protocol, reliability, and sensitivity
Non compliance note. The governance crosswalk indicates plausible evidentiary support, not legal compliance.
Appendix D Glossary (Key Terms & Acronyms)
RAI:
@@ Appendix D Glossary (Key Terms & Acronyms)
Responsible AI: practices and governance to ensure fairness, privacy, safety, transparency, and accountability across the AI lifecycle.
GenAI:
@@
Generative AI: models that synthesize content (text, image, audio, code) from learned distributions.
LLM:
@@
Large Language Model (e.g., GPT, Llama): transformer-based models trained on large text corpora.
VLM:
@@
Vision–Language Model: multimodal models handling text+image inputs/outputs (e.g., BLIP-2, LLaVA).
RAG:
@@
Retrieval-Augmented Generation: generation conditioned on retrieved external context.
Agentic AI:
@@
Systems where models plan, select tools, act, and maintain memory; may be multi-agent.
XAI:
@@
Explainable AI: methods to make model behavior intelligible (ante-hoc/by-design or post-hoc).
LIME:
@@
Local Interpretable Model-agnostic Explanations: local surrogate explanations.
SHAP:
@@
SHapley Additive exPlanations: additive attributions based on cooperative game theory.
Grad-CAM:
@@
Gradient-weighted Class Activation Mapping: visual saliency for CNN/VLM vision backbones.
Calibration:
@@
Agreement between predicted confidence and empirical accuracy.
ECE:
@@
Expected Calibration Error: summary gap between confidence and accuracy across bins.
Brier score:
@@
Mean squared error of probabilistic predictions vs. outcomes (lower is better).
SPD:
Statistical Parity Difference: $\Pr(\hat{y}=1|A=a)-\Pr(\hat{y}=1|A=b)$ (group outcome gap).
DI:
Disparate Impact: $\Pr(\hat{y}=1|A=a)/\Pr(\hat{y}=1|A=b)$ (parity ratio; use $\min(\mathrm{DI},1/\mathrm{DI})$).
Equalized odds:
Parity of TPR/FPR across protected groups (conditional on true label).
Robustness:
Performance stability under noise, distribution shift, or adversarial perturbations.
Red teaming:
Structured adversarial testing to elicit failures or unsafe behavior.
Benchmark:
Standardized tasks/metrics for evaluation (e.g., safety, bias, privacy).
KPI:
Key Performance Indicator: operational metric tracked across the lifecycle (e.g., SPD, ECE, kWh).
DP:
Differential Privacy: formal privacy guarantee bounding information leakage about individuals.
Membership inference:
Attack inferring whether a record was in training data.
Data extraction:
Recovery of memorized training snippets from a model.
PII:
Personally Identifiable Information: data that identifies an individual.
Provenance:
Traceability of data/model/content origin; includes watermarking and content credentials.
Deepfake:
Synthetic media that mimics real persons/events (face/voice/image/video).
System-level failure:
Failures emerging from tool use, planning, memory, or multi-agent interaction.
Governance crosswalk:
Mapping of technical evidence (benchmarks/KPIs) to regulatory obligations.
NIST AI RMF:
U.S. National Institute of Standards and Technology AI Risk Management Framework.
EU AI Act:
European Union regulation with risk tiers and obligations (robustness, transparency, monitoring).
ISO/IEC 42001:
AI management system standard for governance and operational controls.
SAIF:
Secure AI Framework (Google): security practices for AI systems.
AIF360:
IBM AI Fairness 360: fairness metrics and mitigation toolkit.
Fairlearn:
Microsoft toolkit for fairness assessment and mitigation.
Model card:
Standardized documentation of a model’s intended use, data, metrics, and limits.
Datasheet:
Documentation for datasets (provenance, consent/licensing, collection, intended use).
Post-market monitoring:
Ongoing evaluation and incident response after deployment.
Risk families (C1–C10):
Rubric criteria: C1 Bias/Discrimination, C2 Toxicity, C3 Security/Robustness, C4 Mis/Disinformation, C5 Deepfakes/Media Integrity, C6 Privacy, C7 System-level Failure, C8 Adversarial/Red-team Realism, C9 Metric Validity/Documentation, C10 Governance Alignment.
Responsible AI (RAI). A paradigm that embeds fairness, accountability, transparency, privacy, and safety across the AI lifecycle (design, training, deployment, and governance).
Generative AI (GenAI). Models that synthesize novel content (text, image, audio, multimodal) from learned data distributions, including large language models (LLMs), diffusion models, and vision–language models (VLMs).
@@
Bias. Systematic and unfair deviations in model predictions across demographic groups, often stemming from imbalanced or unrepresentative training data.
Fairness. The principle of ensuring equitable outcomes across protected groups, commonly measured through:
$\text{Statistical Parity Difference (SPD)}=P(\hat{y}=1\mid A=a)-P(\hat{y}=1\mid A=b),$
$\text{Disparate Impact (DI)}=\frac{P(\hat{y}=1\mid A=a)}{P(\hat{y}=1\mid A=b)}.$
@@
Hallucination. Model outputs that are factually incorrect, unsupported by evidence, or inconsistent with the input context.
Explainability (XAI). Methods that render black-box model behavior interpretable to humans (e.g., SHAP, LIME, attention maps), enabling accountability and compliance.
Privacy. Safeguarding sensitive data against unauthorized exposure, often operationalized via techniques such as Differential Privacy (DP) and Federated Learning (FL).
Robustness. The ability of models to maintain reliable performance under distribution shifts, adversarial perturbations, or noisy inputs.

### v6

@@ 6.6. Research directions
For practitioners and policy makers, Tables 3 to 8 map benchmarks to risk families and a governance crosswalk provide a starting point for evidence collection. These tables also reveal gaps that are relevant for conformity assessment under emerging regimes. Applied teams can use the matrix to assemble a minimal set of evaluations that cover privacy, robustness, transparency, logging, and post-market monitoring, and then plan additions for multimodal deepfake risk and adaptive red teaming as required by sectoral context.
@@
We highlight five priorities. First, we need to extend benchmarks to capture system-level failures in tool-using and multi-agent settings, especially around retrieval, planning, and actuation (e.g., API calls, code execution, or robotic steps). Second, fold privacy and provenance checks into routine testing, including membership-inference and training-data extraction probes, plus audits of logging sufficiency. Third, create deepfake risk assessments for multimodal models that cover detection, content provenance, and robustness to evasion. Fourth, shift from static to adaptive evaluation with undisclosed items, targeted red teaming, and tests under distribution shift. Fifth, report sustainability alongside accuracy and robustness energy, emissions, and compute-normalized efficiency to enable compute-proportional evaluation and responsible deployment. These priorities directly address the risk-surface gaps surfaced in this survey.
@@
Teams putting ResGenAI into practice can follow a simple loop. First, scope risks by mapping planned features to common risk families (e.g., bias, toxicity, privacy, and security). Next, select baselines from our matrix so that each release is covered by minimum tests across those areas. Then, add context: turn on logging, enforce dataset/model versioning, and keep lightweight documentation so decisions and results remain traceable. Before launch and later during post-market monitorin, there is need to exercise adaptivity with red teaming and a few undisclosed checks to catch regressions. Finally, track a small set of KPIs across the lifecycle: calibration, error rates in high-stakes flows, audit cadence, exclusivity indicators, and energy/emissions. This routine keeps evaluation aligned with governance expectations and makes evidence easier to reuse when audits occur.
6.7. Systemic Vulnerabilities and Blind Spots in ResGenAI
While GenAI capabilities have accelerated, structural blind spots persist across technical and sociotechnical domains (Kenny et al., 2021). This review identifies several systemic gaps that currently hinder universal assurance:
  • RLHF and Alignment Transparency: Alignment advances, such as RLHF and Constitutional AI, have shifted emphasis from raw capability to controllability. However, the documentation of these alignment pathways is often inconsistent, which complicates the auditing of model refusal behaviour and evidence portability across versions (of Standards and, NIST, NIST).
  • Opaque Proprietary Datasets: Responsible stewardship demands adherence to consent and privacy regimes. However, many frontier models rely on web-scale datasets where provenance, licensing, and curation criteria remain undocumented. Truly representative datasets are often infeasible due to data scarcity or legal limits, leading to persistent data leakage and memorization risks (Nasr et al., 2023).
  • Non-Stationary Safety: Finally, safety is inherently non-stationary; model updates, distribution shifts, and evolving adversarial tactics induce drift. Static benchmarks rarely capture these dynamic failure modes , necessitating a shift toward the continuous monitoring and adaptive evaluation frameworks discussed in Section Section 4.2 .
@@ 7. Conclusion
This survey comprehensively synthesized the landscape of ResGenAI by evaluating the alignment between global governance mandates and concrete engineering practices. Our synthesis findings revealed a significant evaluation imbalance in current practices, for example, while technical coverage is mature for surface level AI harms like bias and toxicity, the ecosystem remains acutely vulnerable to privacy leakage, deepfake risks, and the system level failure modes inherent in emergent agentic and tool using environments. Furthermore, we identified that the prevalence of static, task-specific benchmarks severely hinders evidence for institutional audits and regulatory compliance. To bridge this evaluation gap, we proposed a research agenda that prioritizes adaptive, multimodal evaluation and provides a practical stewardship to operationalize technical controls through lifecycle KPIs and standardized audit artifacts. By formalizing the symmetric responsibility model and the ResGenAI audit loop, this work establishes the first integrated framework for bridging technical evaluation with regulatory needs, providing a safe, transparent, and sustainable deployment of GenAI models across domains.
@@
This quote encapsulates the core challenge of designing AI systems that can adapt ethically and responsibly to evolving societal norms. Adaptability in AI must go beyond technical performance to include ethical flexibility, ensuring systems align with human values. ResGenAI emphasizes not just technical capabilities but also the ethical imperatives of transparency, fairness, and accountability.
@@
Achieving ResGenAI requires a collective effort among developers, policymakers, and users to mitigate bias, prioritize privacy, and ensure accountability. Collaboration across disciplines and sectors is essential to create equitable and ethical AI systems that serve diverse societal needs.
@@
Traditional AI benchmarks have prioritized accuracy and efficiency, often neglecting fairness, accountability, and societal impact. Redefining benchmarks to incorporate ethical metrics is critical for guiding the development of AI systems that perform both technically and ethically. The data-driven nature of AI presents persistent challenges, as models often inherit biases and privacy risks from training data. These biases, embedded in AI outputs, require rigorous mitigation strategies to distinguish between useful correlations and harmful patterns. While creating a perfectly unbiased AI system is unattainable due to the evolving nature of societal values, continuous evaluation and adaptation of AI models are essential to maintain alignment with ethical standards.
Human decision-making is influenced by moral reasoning and context, while AI operates strictly within its programmed algorithms. This underscores the importance of fostering AI literacy (Long and Magerko, 2020) among developers and users to ensure responsible deployment and use. To implement ResGenAI effectively, AI systems must prioritize transparency, fairness, and adaptability. They should incorporate ethical reasoning and evolve alongside societal norms. As we advance toward Artificial General Intelligence, the true measure of AI should extend beyond technical proficiency to include ethical adaptability (Blili-Hamelin et al., 2024). The pursuit of ResGenAI is challenging yet indispensable. Ongoing advancements in ResGenAI frameworks, coupled with a growing emphasis on AI ethics, offer a path forward. Prominent voices in the field, including Jensen Huang and Sam Altman, emphasize the need for safe and ethical AI.
@@ 8. Acknowledgement
Appendix A Appendix
A.1. Full Search Strings and Filters
Below we list the precise strings and any database-specific options.
ACM Digital Library Fields: Title/Abstract/Keywords. Years: 2022–2025. Document types: Journal, Proceedings.
@@ A.1. Full Search Strings and Filters
IEEE Xplore Metadata search (Abstract + Index Terms). Years: 2022–2025. Content type: Journals, Conferences.
@@
Web of Science Topic search (TS=Title, Abstract, Keywords). Years: 2022–2025.
@@
arXiv Categories: cs.AI, cs.LG, cs.CV, cs.CL. Date filter applied via API.
@@
Scopus Field search (TITLE-ABS-KEY). Years: 2022–2025.
TITLE-ABS-KEY("Responsible AI" OR "Ethical AI" OR "AI safety" OR transparency OR
accountability OR fairness OR "bias mitigation" OR privacy OR equity OR "open science")
AND TITLE-ABS-KEY("generative AI" OR "large language model*" OR LLM OR LVLM OR
"vision-language" OR multimodal OR "natural language processing" OR "computer vision")
A.2. Codebook for rubric C1 to C10
Use Score only explicit tasks and metrics. If a task maps to multiple risks, assign it to the primary stated objective to avoid double counting. Evidence priority Official paper and docs; then dataset or model cards and maintained leaderboards; then widely cited third party reports.
@@ A.2. Codebook for rubric C1 to C10
Appendix B Scoring protocol, reliability, and sensitivity
Procedure Two reviewers independently scored each criterion using this codebook. Evidence was drawn from official papers, public documentation, and widely cited usage reports, following the priority ordering specified in the codebook. Disagreements were resolved through structured discussion and, where needed, adjudication by a third reviewer. The codebook and scoring definitions are given below:
Reliability Scoring followed a two-reviewer protocol with codebook-guided independent assessment (Appendix A.2) and discussion-based adjudication for disagreements. A third reviewer was consulted when consensus could not be reached. This structured qualitative approach prioritizes transparency and traceability of judgments over formal inter-rater statistics, which is appropriate given the interpretive nature of rubric-based benchmark assessment in survey research. The codebook, scoring definitions, and adjudication procedure are documented to enable replication Indices Report the risk coverage index $S_{\mathrm{cov}}$, coverage rate $R$, assurance index $S_{\mathrm{assur}}$, and the overall normalized index $S$ as defined in Section 4.2.1.
Robustness and Sensitivity Analysis To verify that our conclusions do not depend on arbitrary weighting, we performed a Sensitivity Analysis of the composite scoring index S. While the primary results (Table 4) utilize equal weighting across criteria C1 to C10, we recomputed the scores under alternate weights that prioritized high-stakes risks (e.g., C3 Security and C6 Privacy). We observed a high Spearman rank correlation ($\rho>0.85$) between the equal-weight baseline and the weighted models, indicating that the comparative breadth of benchmarks is a stable indicator of evaluative maturity rather than an artifact of the scoring scheme.
@@ Appendix B Scoring protocol, reliability, and sensitivity
Non compliance note The governance crosswalk indicates plausible evidentiary support, not legal compliance.
B.1. Glossary (Key Terms & Acronyms)
Table 1. Glossary of key terms and acronyms used in this survey.
Term / Acronym Category Concise definition
@@ B.1. Glossary (Key Terms & Acronyms)
RAI Acronym Responsible AI: fairness, privacy, safety, transparency, accountability across the AI lifecycle.
@@
GenAI Acronym Generative AI: models that synthesize content (text/image/audio/code) from learned distributions.
@@
LLM Acronym Large Language Model (e.g., GPT, Llama): transformer-based models trained on large text corpora.
@@
VLM Acronym Vision–Language Model: multimodal models handling text+image inputs/outputs (e.g., BLIP-2, LLaVA).
LMM Acronym Large Multimodal Model: large-scale models spanning multiple modalities (text, vision, audio).
@@
RAG Acronym Retrieval-Augmented Generation: generation conditioned on retrieved external context.
@@
Agentic AI Term Systems where models plan, select tools, act, and maintain memory; may be multi-agent.
CoT Acronym Chain of Thought: intermediate reasoning steps used to solve tasks.
RLHF Acronym Reinforcement Learning from Human Feedback: preference-based alignment fine-tuning.
PEFT Acronym Parameter-Efficient Fine-Tuning: adaptation with limited trainable parameters (e.g., LoRA).
HITL Acronym Human-in-the-Loop: human oversight/feedback integrated into development or operations.
TEVV Acronym Testing, Evaluation, Verification, and Validation: lifecycle assurance practices.
@@
XAI Acronym Explainable AI: methods to make model behavior intelligible (ante-hoc or post-hoc).
@@
LIME Acronym Local Interpretable Model-agnostic Explanations: local surrogate explanations.
@@
SHAP Acronym SHapley Additive exPlanations: additive attributions based on game theory.
@@
Grad-CAM Term Gradient-weighted Class Activation Mapping: visual saliency for CNN/VLM backbones.
@@
Calibration Term Agreement between predicted confidence and empirical accuracy.
@@
ECE Acronym Expected Calibration Error: bin-wise gap between confidence and accuracy (summary).
@@
Brier score Term Mean squared error of probabilistic predictions vs. outcomes (lower is better).
@@
Bias Term Systematic and unfair deviations across demographic groups, often from unrepresentative data.
Fairness Term Equitable outcomes across protected groups; often assessed via SPD/DI and related metrics.
Equalized odds Term Parity of TPR/FPR across protected groups (conditional on true label).
Robustness Term Performance stability under noise, distribution shift, or adversarial perturbations.
Red teaming Term Structured adversarial testing to elicit failures or unsafe behavior.
Benchmark Term Standardized tasks/metrics for evaluation (e.g., safety, bias, privacy).
KPI Acronym Key Performance Indicator: operational metric tracked across the lifecycle (e.g., SPD, ECE, kWh).
@@
Hallucination Term Outputs that are incorrect, unsupported by evidence, or inconsistent with input/context.